OpenStudy (anonymous):

Let f(x) =9x - 4.5 Find the value of f(2) Answers: A. 6.5 B. 87.5 C. 13.5 D. 9

OpenStudy (texaschic101):

just sub in 2 for x f(2) = 9(2) - 4.5 f(2) = 18 - 4.5 can you do the math and finish the problem ?

OpenStudy (anonymous):

Thanks !!!! I got it :)